A gun with inertia 4.5 kg fires a 12-g bullet at a stationary target located 1.0 km away. After the bullet leaves the gun, its speed decreases (constant x component of acceleration a = -1.0 m/s2 ) so that the bullet hits the target at 299 m/s.
If the direction in which the bullet moves is along a positive x coordinate axis, what is the recoil velocity of the gun?
1
Expert's answer
2020-04-28T09:26:57-0400

As per the given question,

Mass of the gun"(M)=4.5 kg"

Mass of the bullet"(m)=12gm=12\times 10^{-3}kg"

Distance of the target "(d)=1k =1000 m"

deceleration of the bullets "a=-1 m\/sec^2"

Hitting speed of the bullet to the target "(v)=299 m\/sec"

Now, Let the initial speed of the bullet =u

So,

"v^2=u^2-2ad"

"\Rightarrow u^2=v^2+2ad"

"\Rightarrow v=\sqrt{299^2+2\times 1\times 1000}=302.32 m/sec"

Now applying the conservation of momentum,

Let the recoil speed of the gun "v_1"

"Mv_1+mu=0"

"\Rightarrow v_2=\dfrac{-mu}{ M}=\dfrac{-12\times 10^{-3}\times 302.32}{4.5}=-0.806 m/sec"

Here negative sign is representing that the bullet will move to negative x axis with the mentioned speed.